# Radar Perception in Autonomous Driving: Exploring Different Data Representations

### v1

@@ I Introduction
Nowadays, the automotive industry has witnessed significant advancements in autonomous driving technologies, revolutionizing intelligent vehicles and intelligent transportation. With the aim of creating safer and more efficient roadways, one crucial aspect is the development of reliable sensor perception for autonomous driving systems. Compared to LiDARs and cameras, the fundamental advantage of radar sensors lies in their capabilities in range, velocity and angle measurements [1, 2]. Additionally, radars have robust sensing capabilities, as exhibited superior effectiveness under adverse lighting and weather conditions [3, 4]. Moreover, radar sensors can even perceive objects behind walls, providing vehicles with early warning of potential obstacles or hazards [5, 6]. All of the above characteristics make radar an indispensable component in ensuring the reliability and superiority of autonomous driving systems, especially in scenarios where other sensor inputs might be limited or compromised.
@@
  • We offer an up-to-date (2019-2023) overview of radar-based datasets and algorithms, providing in-depth research on their advantages and limitations.
@@
The rest of this review is structured as follows: Section II provides an overview of radar technology, including its working principles and signal processing techniques employed in radar perception. Section III presents an in-depth examination on datasets and methods of different radar data representations. Section IV discusses the challenges and potential directions for research and development for radar perception in autonomous driving. Lastly, Section V summarizes our study and presents an outlook for future works.
@@ II-B1 Parameter Estimation
Range
${R}=\frac{cf_{b}}{2S}$ (1)
${R_{res}}=\frac{c}{2B}$ (2)
${c}$ is the speed of the light ($3*10^{8}m/s$). ${f_{b}}$ is the instantaneous frequency difference at the mixers from TX-chirp and RX-chirp. ${S}$ is the slope of a chirp. ${B}$ refers to the frequency variation range.
Doppler Velocity
${V}=\frac{\lambda\Delta\phi}{4\pi T_{c}}$ (3)
${V_{res}}=\frac{\lambda}{2T_{f}}$ (4)
$\lambda$ is the wavelength of the light. ${\Delta\phi}$ is a shift in phase. ${T_{c}}$ is the duration between the chirps. ${T_{f}}$ is cycle time of a frame.
Azimuth Angle
$\theta=sin^{-1}(\frac{\lambda\Delta\phi}{2\pi l_{a}})$ (5)
$\theta_{res}=\frac{\lambda}{N_{a}l_{a}cos(\theta)}$ (6)
$l_{a}$ is the length between RX antennas in azimuth. $N_{a}$ is the number of RX antennas in azimuth.
Elevation Angle
$\theta=sin^{-1}(\frac{\lambda\Delta\phi}{2\pi l_{e}})$ (7)
$\theta_{res}=\frac{\lambda}{N_{e}l_{e}cos(\theta)}$ (8)
$l_{e}$ is the length between RX antennas in elevation. $N_{e}$ is the number of RX antennas in elevation.
@@ III Radar Data Representations
Fig. 5 illustrates the generation progress of the five different radar data representations. Initially, the raw output from the radar sensor, known as the ADC signal, is hardly interpretable by human observers. Thus, researchers utilize 3D FFTs across the sample, chirp, and antenna dimensions to transform the ADC signal into an image-like representation called the radar tensor. Subsequently, peak detection is employed on the radar tensor to eliminate clutter, resulting in a sparse, point-like representation referred to as the point cloud. By accumulating point clouds over a specific duration and applying grid mapping methods, the grid map representation is generated for the purpose of identifying static objects. Moreover, some researchers perform a Time-Frequency transform following the Range-FFT to extract the micro-Doppler signature, which is utilized for recognizing objects characterized by micro-motion features.
@@
RaDICaL [38]
ADC Signal
2021
Object Detection
Radar (TI IWR1443), RGB-D Camera
@@
Indoor (people, static clutter), Roads (urban, rural, highway, various traffic scenarios)
RADIal [39]
ADC Signal, Radar Tensor, Point Cloud
2021
Object Detection, Semantic Segmentation
Radar (high-definition), Cameras, LiDAR
Roads (urban, highway, rural)
CARRADA [40]
Radar Tensor
2020
Detection, Semantic Segmentation, Object Tracking
Radar (TI AWR1843), RGB-D Camera, LiDAR
Roads (urban, highway, intersection scenarios)
Zendar [41]
Radar Tensor, Point Cloud
2020
Object Detection, Mapping, Localization
Radar (synthetic aperture), Camera, LiDAR
Roads (diverse urban driving environments)
RADIATE [42]
Radar Tensor
2020
Object Detection
Radar (Navtech CTS350-X), Camera
@@
Roads (wet, snowy, foggy, rainy, nighttime, urban, highway)
MulRan [43]
Radar Tensor
2020
Place Recognition
Radar (Navtech CIR204-H), Cameras, LiDAR
@@
Roads (city, highway, intersection, crosswalks, parks, recreational areas, tunnels, bridges)
Oxford Radar RobotCar [44]
Radar Tensor, Grid Map
2020
Object Detection, Odometry
Radar (Navtech CTS350-X), Camera, LiDAR, GPS, INS
@@
Roads (urban, highway, rural, industrial area, residential area, roundabout, intersection)
SCORP [45]
Radar Tensor
2020
Semantic Segmentation
Radar (76 GHz), Camera
Roads (parking lot)
CRUW [46]
Radar Tensor
2021
Object Detection
Radar (TI AWR1843, DCA1000), Cameras
@@
Roads (parking, campus, city, highway)
RADDet [47]
Radar Tensor
2021
Object Detection
Radar (TI AWR1843), Stereo Cameras
@@
Roads (urban, rural, highway, intersections, weather conditions)
Boreas [48]
Radar Tensor
2022
Object Detection, Localization, Odometry
Radar (Navtech CIR304-H), Camera, LiDAR
Roads (highway, rural, urban)
ColoRadar [49]
@@
Radar Tensor, Point Cloud
2022
Localization
Radar (AWR1843), LiDAR, IMU
Indoor, outdoor environments
K-Radar [50]
Radar Tensor
2022
Object Detection, Object Tracking, SLAM
Radar (RETINA-4ST), Stereo Cameras, LiDAR
Roads (highway, intersection, urban)
nuScenes [51]
Point Cloud, Grid Map
2019
Object Detection, Object Tracking
Radar (Continental ARS408), Camera, LiDAR
Roads (intersection, crosswalk, roundabout, pedestrian crossing)
Astyx [52]
Point Cloud
2019
Object Detection
Radar (Astyx 6455 HiRes), Camera, LiDAR
@@
Roads (highway, urban, rural, parking, roundabout)
SeeingThroughFog [53]
Point Cloud
2020
Object Detection
Radar (77GHz), Stereo/Gated/FIR Cameras, LiDAR
@@
Adverse road conditions (clear, rainy, snowy, foggy, nighttime, urban, highway, rural, traffic)
AIODrive [54]
Point Cloud
2020
Object Detection, Semantic Segmentation, Object Tracking, Depth Estimation
Radar (77GHz), RGB/Stereo Cameras, LiDAR
Roads (highway, residential street, parking)
RadarScenes [55]
Point Cloud
2021
Semantic Segmentation, Object Tracking
Radar (77GHz), Documentary Camera
@@
Roads (urban, suburban, rural, highway, tunnel, intersection, roundabout, parking)
Pixset [56]
Point Cloud
2021
Object Detection, Object Tracking
Radar (TI AWR1843), Cameras, LiDARs
Roads (urban, suburban, highway)
VoD [57]
Point Cloud
2022
Object Detection, Object Tracking
Radar (ZF FRGen 21), Stereo Camera, LiDAR
Roads (highway, rural, urban)
TJ4DRadSet [58]
Point Cloud
2022
Object Detection, Object Tracking
Radar (Oculii Eagle), Camera, LiDAR
Roads (intersections, one-way streets)
aiMotive [59]
Point Cloud
2022
Object Detection
Radar (77GHz), Camera, LiDAR, GPS, IMU
Roads (highway, urban, rural)
WaterScenes [60]
Point Cloud
2023
Object Detection, Instance Segmentation,
Radar (Oculii Eagle), Camera, GPS, IMU
Waterways (river, lake, canal, moat)
NTU4DRadLM [61]
Point Cloud
2023
SLAM
@@
Radar (Oculii Eagle), RGB/Thermal Cameras, LiDAR, GPS, IMU
Roads (carpark, garden, campus)
Dual-Radar [62]
Point Cloud
2023
Object Detection, Object Tracking
Radar (ARS548 RDI, Arbe Phoenix), Camera, LiDAR
Roads (urban, suburban, highway, tunnel, parking)
Dop-NET [63]
Micro-Doppler Signature
2020
Classification
Radar (Ancortek 24GHz)
@@
Gestures (wave, pinch, click, swipe)
CI4R [64]
Micro-Doppler Signature
2020
Classification
Radar (77GHz, 24GHz, Xethru)
@@
Activities (walking, picking, sitting, crawling, kneeling, limping)
Open Radar Datasets [65]
Micro-Doppler Signature
2021
Classification
Radar (TI AWR2243), Camera, GPS, IMU
Roads (urban, highway, rural)
MCD-Gesture [66]
Micro-Doppler Signature
2022
Classification
Radar (TI AWR1843)
@@
Gestures (push, pull, slide left, slide right, clockwise turning, counterclockwise turning)
@@ III-A ADC Signal
When an analog signal is sampled and quantized by an ADC, it produces a sequential data stream called the ADC signal. As the raw data produced by radar sensors, ADC signals retain all information from the detections, which is highly valuable in deep learning applications. At this stage, the signal lacks spatial coherence among its values, as all information is confined to the time domain [45]. To represent the ADC signal in a more structured format, it is common to transform it into a 3D Sample-Chirp-Antenna (SCA) tensor, as illustrated in Fig. 3(e).
@@ III-A1 Datasets
With the development of radar technology and increased computing capability, radar data characterized by ADC signal has emerged and garnered widespread attention in recent years. RaDICaL [38] is the first dataset that offers raw ADC signal data collected explicitly for road scenarios in autonomous driving. With access to raw radar measurements, the authors encouraged researchers to design novel processing methods to perform object detection directly or to get downstream data representations. RADIal [39] is the most comprehensive dataset in terms of radar data representations. In addition to providing ADC signals, the RADIal dataset contains processed data derived from the ADC signals, including radar tensors and point clouds. Apart from using ADC signals in road scenarios, another significant application is in-cabin vital signs monitoring. However, existing datasets in this area (e.g., [67, 68, 69, 70]) primarily focus on indoor scenarios like hospitals or home environments, which differ from in-cabin scenarios in autonomous driving. Although some studies (e.g., [9, 10, 71]) have investigated monitoring occupant vital signs in cabins, their datasets are not publicly available.
@@ III-A2 Methods
In recent years, the development of road object perception using ADC signals has gradually gained momentum, highlighted in motion classification (e.g., [72, 73, 74, 75]) and object detection (e.g., [76, 77, 78, 79]). These studies broadly encompass two research directions. Primarily, ADCNet [76], T-FFTRadNet [77], and CubeLearn [75] involve designing individual end-to-end learnable radar perception models, enabling object perception on ADC signals. These approaches utilize deep learning on ADC signals as a substitute for traditional signal processing procedures. This solution replaces the computationally intensive FFTs and simplifies the data flow in embedded implementations, thus significantly reducing the computational requirements.
Another direction of road object-based methods focuses on improving the resolution of radar data, which is essential for estimating object locations and velocities. For example, the ADC Super-Resolution (ADC-SR) model [79] leverages ADC signals to enhance the radar’s azimuth resolution, tackling complexities in ADC signals and predicting signals from unseen receivers. As such, the hallucinated ADC signals can further refine RAD tensors to enhance object detection capabilities in autonomous driving.
@@
The vital sign-based methods focus on monitoring the physiological signals of occupants inside the vehicle cabin, including heart rate, respiration rate, and presence detection [80, 9, 71, 81, 82]. Signal processing techniques, such as spectrogram analysis, short-time Fourier transform or wavelet transform, are employed to extract the vital sign signals from the radar ADC signal [8, 82]. Machine learning algorithms, such as statistical classifiers or regression models, may be trained using labeled data to estimate vital sign parameters from the extracted features [8].
Radar systems used for cabin vital sign monitoring are generally designed for close-range operation, such as within a vehicle or confined space. They focus on high-resolution detection of small movements and physiological changes, and may use less power and have smaller form factors than road-oriented radar systems.

### v2

# Exploring Radar Data Representations in Autonomous Driving: A Comprehensive Review

@@ I Introduction
Nowadays, the automotive industry has witnessed significant advancements in autonomous driving technologies, revolutionizing intelligent vehicles and intelligent transportation. With the aim of creating safer and more efficient roadways, one crucial aspect is the development of reliable sensor perception for autonomous driving systems. Compared to LiDARs and cameras, the fundamental advantage of radar sensors lies in their capabilities in range, velocity and angle measurements [1, 2]. Additionally, radars have robust sensing capabilities and exhibit superior effectiveness under adverse lighting and weather conditions [3, 4]. Moreover, radar sensors can even perceive objects behind walls, providing vehicles with early warning of potential obstacles or hazards [5, 6]. All of the above characteristics make radar an indispensable component in ensuring the reliability and superiority of autonomous driving systems, especially in scenarios where other sensor inputs might be limited or compromised.
@@
Figure 2: Overview of this review. Section II provides an overview of radar perception, including its working principles and signal processing techniques. Section III presents an in-depth examination on datasets and methods of different radar data representations. Section IV discusses the challenges and potential directions for research and development for radar perception in autonomous driving.
@@
  • We offer an up-to-date (2019-2024) overview of radar-based datasets and algorithms, providing in-depth research on their advantages and limitations.
@@
The remainder of this review is structured in Fig. 2 and is described as follows: Section II provides an overview of radar working principles and signal processing techniques, serving as the foundation for generating various radar data representations. Section III presents an in-depth examination on datasets and methods for each radar data representation. Additionally, this section discusses the advantages and limitations of each radar data representation, assisting readers in selecting the most suitable approach for their research. Section IV delves into challenges and potential directions in research and development for radar perception in autonomous driving, guiding readers to focus on current hot topics and explore feasible solutions. Lastly, Section V summarizes our study and presents an outlook for future works, inspiring researchers to make further strides in radar perception.
@@ II-B1 Parameter Estimation
Range ${R}=\frac{cf_{b}}{2S}$ (1) ${R_{res}}=\frac{c}{2B}$ (2) ${c}$ is the speed of the light ($3*10^{8}m/s$). ${f_{b}}$ is the instantaneous frequency difference at the mixers from TX-chirp and RX-chirp. ${S}$ is the slope of a chirp. ${B}$ refers to the frequency variation range.
Doppler Velocity ${V}=\frac{\lambda\Delta\phi}{4\pi T_{c}}$ (3) ${V_{res}}=\frac{\lambda}{2T_{f}}$ (4) $\lambda$ is the wavelength of the light. ${\Delta\phi}$ is a shift in phase. ${T_{c}}$ is the duration between the chirps. ${T_{f}}$ is cycle time of a frame.
Azimuth Angle $\theta=sin^{-1}(\frac{\lambda\Delta\phi}{2\pi l_{a}})$ (5) $\theta_{res}=\frac{\lambda}{N_{a}l_{a}cos(\theta)}$ (6) $l_{a}$ is the length between RX antennas in azimuth. $N_{a}$ is the number of RX antennas in azimuth.
Elevation Angle $\theta=sin^{-1}(\frac{\lambda\Delta\phi}{2\pi l_{e}})$ (7) $\theta_{res}=\frac{\lambda}{N_{e}l_{e}cos(\theta)}$ (8) $l_{e}$ is the length between RX antennas in elevation. $N_{e}$ is the number of RX antennas in elevation.
@@ III Radar Data Representations
Fig. 6 illustrates the generation process of five different radar data representations. Initially, the raw output from the radar sensor, known as the ADC signal, is hardly interpretable by human observers. Thus, researchers utilize 3D FFTs across the sample, chirp, and antenna dimensions to transform the ADC signal into an image-like representation called the radar tensor. Subsequently, peak detection is employed on the radar tensor to eliminate clutter, resulting in a sparse, point-like representation referred to as the point cloud. By accumulating point clouds over a specific duration and applying grid mapping methods, the grid map representation is generated for the purpose of identifying static objects. Moreover, some researchers perform a Time-Frequency transform following the Range-FFT to extract the micro-Doppler signature, which is utilized for recognizing objects characterized by micro-motion features.
@@
RaDICaL [38] ADC Signal 2021 Object Detection Radar (TI IWR1443), RGB-D Camera Indoor (people, static clutter), Roads (urban, rural, highway, various traffic scenarios)
RADIal [39] ADC Signal, Radar Tensor, Point Cloud 2021 Object Detection, Semantic Segmentation Radar (high-definition), Cameras, LiDAR Roads (urban, highway, rural)
CARRADA [40] Radar Tensor 2020 Detection, Semantic Segmentation, Object Tracking Radar (TI AWR1843), RGB-D Camera, LiDAR Roads (urban, highway, intersection scenarios)
Zendar [41] Radar Tensor, Point Cloud 2020 Object Detection, Mapping, Localization Radar (synthetic aperture), Camera, LiDAR Roads (diverse urban driving environments)
@@
RADIATE [42] Radar Tensor 2020 Object Detection Radar (Navtech CTS350-X), Camera Roads (wet, snowy, foggy, rainy, nighttime, urban, highway)
@@
MulRan [43] Radar Tensor 2020 Place Recognition Radar (Navtech CIR204-H), Cameras, LiDAR Roads (city, highway, intersection, crosswalks, parks, recreational areas, tunnels, bridges)
@@
Oxford Radar RobotCar [44] Radar Tensor, Grid Map 2020 Object Detection, Odometry Radar (Navtech CTS350-X), Camera, LiDAR, GPS, INS Roads (urban, highway, rural, industrial area, residential area, roundabout, intersection)
SCORP [45] Radar Tensor 2020 Semantic Segmentation Radar (76 GHz), Camera Roads (parking lot)
@@
CRUW [46] Radar Tensor 2021 Object Detection Radar (TI AWR1843, DCA1000), Cameras Roads (parking, campus, city, highway)
@@
RADDet [47] Radar Tensor 2021 Object Detection Radar (TI AWR1843), Stereo Cameras Roads (urban, rural, highway, intersections, weather conditions)
Boreas [48] Radar Tensor 2022 Object Detection, Localization, Odometry Radar (Navtech CIR304-H), Camera, LiDAR Roads (highway, rural, urban)
@@
ColoRadar [49] Radar Tensor, Point Cloud 2022 Localization Radar (AWR1843), LiDAR, IMU Indoor, outdoor environments
K-Radar [50] Radar Tensor 2022 Object Detection, Object Tracking, SLAM Radar (RETINA-4ST), Stereo Cameras, LiDAR Roads (highway, intersection, urban)
OORD [51] Radar Tensor 2024 Place Recognition Radar (Navtech CTS350-X), GPS/INS Roads (off-road, difficult terrain, naturalistic environments)
nuScenes [52] Point Cloud, Grid Map 2019 Object Detection, Object Tracking Radar (Continental ARS408), Camera, LiDAR Roads (intersection, crosswalk, roundabout, pedestrian crossing)
@@
Astyx [53] Point Cloud 2019 Object Detection Radar (Astyx 6455 HiRes), Camera, LiDAR Roads (highway, urban, rural, parking, roundabout)
@@
SeeingThroughFog [54] Point Cloud 2020 Object Detection Radar (77GHz), Stereo/Gated/FIR Cameras, LiDAR Adverse road conditions (clear, rainy, snowy, foggy, nighttime, urban, highway, rural, traffic)
AIODrive [55] Point Cloud 2020 Object Detection, Semantic Segmentation, Object Tracking, Depth Estimation Radar (77GHz), RGB/Stereo Cameras, LiDAR Roads (highway, residential street, parking)
@@
RadarScenes [56] Point Cloud 2021 Semantic Segmentation, Object Tracking Radar (77GHz), Documentary Camera Roads (urban, suburban, rural, highway, tunnel, intersection, roundabout, parking)
Pixset [57] Point Cloud 2021 Object Detection, Object Tracking Radar (TI AWR1843), Cameras, LiDARs Roads (urban, suburban, highway)
VoD [58] Point Cloud 2022 Object Detection, Object Tracking Radar (ZF FRGen 21), Stereo Camera, LiDAR Roads (highway, rural, urban)
TJ4DRadSet [59] Point Cloud 2022 Object Detection, Object Tracking Radar (Oculii Eagle), Camera, LiDAR Roads (intersections, one-way streets)
aiMotive [60] Point Cloud 2022 Object Detection Radar (77GHz), Camera, LiDAR, GPS, IMU Roads (highway, urban, rural)
WaterScenes [61] Point Cloud 2023 Object Detection, Instance/Semantic/Waterline/ Panoptic Segmentation Radar (Oculii Eagle), Camera, GPS, IMU Waterways (river, lake, canal, moat)
@@
NTU4DRadLM [62] Point Cloud 2023 SLAM Radar (Oculii Eagle), RGB/Thermal Cameras, LiDAR, GPS, IMU Roads (carpark, garden, campus)
Dual-Radar [63] Point Cloud 2023 Object Detection, Object Tracking Radar (ARS548 RDI, Arbe Phoenix), Camera, LiDAR Roads (urban, suburban, highway, tunnel, parking)
MiliPoint [64] Point Cloud 2024 Activity Recognition Radar (TI IWR1843), Stereo Camera Activities (identification, action classification and keypoint estimation)
@@
Dop-NET [65] Micro-Doppler Signature 2020 Classification Radar (Ancortek 24GHz) Gestures (wave, pinch, click, swipe)
@@
CI4R [66] Micro-Doppler Signature 2020 Classification Radar (77GHz, 24GHz, Xethru) Activities (walking, picking, sitting, crawling, kneeling, limping)
Open Radar Datasets [67] Micro-Doppler Signature 2021 Classification Radar (TI AWR2243), Camera, GPS, IMU Roads (urban, highway, rural)
@@
MCD-Gesture [68] Micro-Doppler Signature 2022 Classification Radar (TI AWR1843) Gestures (push, pull, slide left, slide right, clockwise turning, counterclockwise turning)
@@ III-A ADC Signal
When an analog signal is sampled and quantized by an ADC, it produces a sequential data stream called the ADC signal. As the raw data produced by radar sensors, ADC signals retain all information from the detections, which is highly valuable for deep learning applications. At this stage, the signal lacks spatial coherence among its values, as all information is confined to the time domain [45]. To be represented in a more structured format, the ADC signal is typically converted to a 3D Sample-Chirp-Antenna (SCA) tensor, as illustrated in Fig. 4(e).
@@ III-A1 Datasets
With the development of radar technology and increased computing capability, radar data characterized by ADC signal has emerged and garnered widespread attention in recent years. RaDICaL [38] is the first dataset that offers raw ADC signal data collected explicitly for road scenarios in autonomous driving. With access to raw radar measurements, the authors encouraged researchers to design novel processing methods to perform object detection directly or to get downstream data representations. RADIal [39] is the most comprehensive dataset in terms of radar data representations. In addition to providing ADC signals, the RADIal dataset contains processed data derived from the ADC signals, including radar tensors and point clouds. Apart from using ADC signals in road scenarios, another significant application is in-cabin vital sign monitoring. However, existing datasets in this area (e.g., [69, 70, 71, 72]) primarily focus on indoor scenarios like hospitals or home environments, which differ from in-cabin scenarios in autonomous driving. Although some studies (e.g., [9, 10, 73]) investigated monitoring occupant vital signs in cabins, their datasets are unavailable publicly.
@@ III-A2 Methods
In recent years, there has been a growing interest in the development of road object perception using ADC signals, as evidenced by research in motion classification (e.g., [74, 75, 76, 77]) and object detection (e.g., [78, 79, 80, 81]). These studies encompass two main research directions. The first direction involves designing individual end-to-end learnable radar perception models (e.g., ADCNet [78], T-FFTRadNet [79], and CubeLearn [77]), enabling object perception on ADC signals. These approaches utilize deep learning on ADC signals as an alternative to traditional signal processing procedures. This solution replaces computationally intensive FFTs and simplifies data flow in embedded implementations, thus significantly reducing the computational requirements.
Another direction of road object-based methods focuses on improving the resolution of radar data, which is essential for estimating object locations and velocities. For example, the ADC Super-Resolution (ADC-SR) model [81] leverages ADC signals to enhance the radar’s azimuth resolution, tackling complexities in ADC signals and predicting signals from unseen receivers. Consequently, the hallucinated ADC signals can further refine RAD tensors, enhancing object detection capabilities in autonomous driving.
@@
The vital sign-based methods focus on monitoring the physiological signals of occupants inside the vehicle cabin, including heart rate, respiration rate, and presence detection [82, 9, 73, 83, 84]. Signal processing techniques, such as spectrogram analysis, Short Time Fourier Transform (STFT) and Wavelet Transform (WT), are employed to extract the vital sign signals from the radar ADC signal [8, 84]. Machine learning algorithms, such as statistical classifiers or regression models, may be trained using labeled data to estimate vital sign parameters from the extracted features [8].
Radar systems used for cabin vital sign monitoring are generally designed for close-range operation, such as within a vehicle or confined space. They focus on high-resolution detection of small movements and physiological changes, and may use less power and have smaller form factors than road-oriented radar systems. However, vital sign-based methods using machine learning algorithms may have limitations in real-time processing, which is crucial in low-latency applications [85, 86].